OpenStudy (anonymous):

Solve and check your answer 2/1-x= 5/1+x

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

\[\frac{2}{1-x}=\frac{5}{1+x}\iff 2(1+x)=5(1-x)\]

OpenStudy (anonymous):

holy what. your so quick!

OpenStudy (saifoo.khan):

He's Chuck Norris! D:

OpenStudy (anonymous):

clearly!

OpenStudy (anonymous):

wait so would you set those to zero to get your solution set?

OpenStudy (saifoo.khan):

Nope. Solve for x.

OpenStudy (anonymous):

So -1, 1

OpenStudy (anonymous):

??

OpenStudy (anonymous):